OpenStudy (meilendurcer):

Which ratios form a proportion? Use cross products to test each pair. A. 15/40, 35/90 B. 18/22, 45/55 C. 15/55, 33/121 D. 17/31, 32/45

OpenStudy (anonymous):

15/55,33/121

OpenStudy (meilendurcer):

Which ratios form a proportion? Use cross products to test each pair. A. 9/17, 25/50 B. 12/35, 24/75 C. 15/55, 33/121 D. 12/35, 20/70

OpenStudy (anonymous):

15/55=3/11 33/121=3/11.....hence they are in proportion

OpenStudy (meilendurcer):

oh... ok

OpenStudy (anonymous):

test for each pair....no othr is satisfying the condition

OpenStudy (meilendurcer):

ok

OpenStudy (meilendurcer):

Which ratios do not form a proportion? A. 24/96, 27/108 B. 17/102, 12/70 C. 21/51, 35/85 D. 45/75, 63/105

OpenStudy (anonymous):

17/102=1/6 12/70=6/35....hence they dont form proportion